kindly explain the answer.
Block A on the left has mass 1.00kg. Block B on the right has mass 3.00kg.The blocks are forced together, compressing the spring.Then the system is released from rest on a level,frictionless surface. After the blocks are released, the kinetic energy (KE) of block A is?
1
Expert's answer
2020-08-10T19:53:15-0400

From the conservation of momentum:


"p_A=p_B\\K_A=\frac{p_A^2}{2m_A}=\frac{p_B^2m_B}{2m_Am_B}=\frac{m_B}{m_A}K_B\\K_A=\frac{3}{1}K_B=3K_B"